### v1

# LLM Agents for Education: Advances and Applications

Abstract.
Large Language Model (LLM) agents have demonstrated remarkable capabilities in automating tasks and driving innovation across diverse educational applications. In this survey, we provide a systematic review of state-of-the-art research on LLM agents in education, categorizing them into two broad classes: (1) Pedagogical Agents, which focus on automating complex pedagogical tasks to support both teachers and students; and (2) Domain-Specific Educational Agents, which are tailored for specialized fields such as science education, language learning, and professional development. We comprehensively examine the technological advancements underlying these LLM agents, including key datasets, benchmarks, and algorithmic frameworks that drive their effectiveness. Furthermore, we discuss critical challenges such as privacy, bias and fairness concerns, hallucination mitigation, and integration with existing educational ecosystems. This survey aims to provide a comprehensive technological overview of LLM agents for education, fostering further research and collaboration to enhance their impact for the greater good of learners and educators alike.
Large language models, Adaptive learning, AI for education
*Equal contribution. Corresponding author.
copyright: nonejournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
1. Introduction
Artificial intelligence (AI) techniques are increasingly used in education to enable personalized learning and intelligent tutoring (Chen et al., 2020; Zhai et al., 2021; Pedro et al., 2019). While traditional educational data mining approaches (Shafiq et al., 2022; Khan and Ghosh, 2021; Abdelrahman et al., 2023; Song et al., 2022; Wang et al., 2022; Gao et al., 2021), such as knowledge tracing and cognitive diagnosis, have made significant progress in reshaping the human-learning paradigm by analyzing student behaviors and assessing knowledge states, they still face major challenges in real-world applications. These challenges include shallow contextual understanding, limited interactive capabilities, and difficulties in generating adaptive, personalized learning materials, etc (Zhu et al., 2024; Laak and Aru, 2024; Tan et al., 2023).
The strong natural language understanding of Large Language Models (LLMs) and the task automation capabilities of LLM agents make them valuable for addressing challenges in education (Weng, 2023; Wang et al., 2024d). First, memory enables LLM agents to retain both long-term knowledge about students’ study habits and short-term context from real-time interactions, enhancing contextual understanding and ensuring personalized learning experiences across various educational tasks (Zhang et al., 2024a). Second, tool use allows LLM agents to access external resources, perform complex calculations, and retrieve real-time information, enabling them to automate intricate educational tasks such as grading, knowledge retrieval, and adaptive content generation, thereby overcoming limited interactivity and enhancing engagement (Gao et al., 2023). Third, planning supports structured learning by decomposing complex topics, predicting optimal learning paths, and dynamically adjusting instructional strategies, allowing LLM agents to autonomously guide students through personalized learning experiences (Huang et al., 2024a). By integrating these capabilities, LLM agents not only enhance understanding, engagement, and personalization but also automate complex educational workflows, making learning more adaptive, efficient, and scalable.
In this survey, we provide a comprehensive review of LLM agents in education. We begin by introducing LLM agents, followed by a general discussion of their potential applications in educational systems. Next, we categorize them into two broad classes: (1) Pedagogical Agents, which automate complex pedagogical tasks such as classroom simulation and learning resource recommendation for teachers. Also, they transform student support paradigms through applications like knowledge tracing, adaptive learning, and error detection and correction. (2) Domain-Specific Educational Agents, which are designed to address specialized challenges in different fields, including science education, language learning, and professional development. Furthermore, we discuss the key challenges in LLM agents for education, which need to be addressed to advance research and enable effective deployment in real-world educational applications. Finally, we compile comprehensive datasets, benchmarks, and evaluation methodologies to support and encourage further research involvement in the area of LLM agents for education. We summarize our contributions as follows:
Figure 1. The overview of LLM Agents for education.
  • Comprehensive review of LLM Agents in education. We analyze recent advancements in LLM-driven educational agents, discussing their applications, effectiveness, and challenges in automating pedagogical tasks and enhancing learning experiences.
  • Novel task-centric taxonomy. We propose a structured classification of LLM agents, categorizing them into Pedagogical Agents and Domain-Specific Educational Agents, providing a framework for understanding their roles and capabilities.
  • Current challenges and future research directions. We analyze critical challenges that need to be addressed for the effective deployment of LLM agents in education, including issues related to privacy, bias and fairness concerns, hallucination, and integration into real-world educational ecosystems.
  • Compilation of essential resources. We compile comprehensive datasets and benchmarks to support future research efforts and facilitate the development of more robust and effective LLM-driven educational solutions [^1] Due to the page limit, we present more details in Appendix A. .
2. LLM Agents for Education
LLM agents, powered by LLMs (OpenAI, 2022, 2023; Touvron et al., 2023), combine core components like memory, tool use, and planning to enhance the performance for diverse tasks (Xi et al., 2025; Durante et al., 2024; Tran et al., 2025; Liu et al., 2025; Yan et al., 2024d). They leverage long-term memory for foundational knowledge and short-term memory for real-time adaptation, enabling context-aware interactions. Active memory management, such as summarization and retrieval (Chen et al., 2023; Liang et al., 2023), helps maintain relevant information. To overcome LLM limitations, such as knowledge cutoff and calculation inefficiency (Liska et al., 2022), LLM agents integrate external tools like search engines, calculators, and databases (Qin et al., 2023; Zhuang et al., 2023), enabling real-time data access and multi-step tasks. Additionally, planning capabilities allow agents to decompose goals, simulate state transitions, and optimize actions for complex problem-solving (Yao et al., 2023; Valmeekam et al., 2023; Xie et al., 2024c).
Education, a field inherently focused on individualized learning and problem-solving, is one of the ideal scenarios for deploying LLM agents. These agents can significantly enhance the learning experience by providing personalized support tailored to the unique needs of each student. To achieve this, as illustrated in Figure 1, LLM agents must integrate several key capabilities, including memory, tool-use, and planning modules. The memory module enables the agent to retain an understanding of fundamental concepts, track the student’s past interactions and progress, and store personalized profiles to provide accurate responses. The tool-use module allows the agent to access specialized educational tools, such as knowledge tracing and recommendation models, as well as external resources like search engines, thereby expanding its functional capabilities and ensuring up-to-date, accurate information. The planning module empowers the agent to deconstruct the student’s objectives, plan tailored learning paths, and offer personalized feedback based on the student’s progress and performance. Together, these modules allow LLM agents to function as dynamic, responsive tutors capable of guiding students through both simple and complex tasks. By offering context-aware feedback, adapting to evolving needs, and tailoring learning experiences, LLM agents have the potential to revolutionize education. Their ability to provide scalable, adaptive, and efficient support can transform personalized education, making it more accessible, engaging, and effective. As shown in Figure 2, we categorize LLM agents for education into two types: Pedagogical Agents and Domain-Specific Educational Agents, and outline the roadmap for discussion in the following sections.
{forest}
for tree= forked edges, grow’=0, calign=child edge, calign child=(n_children()+1)/2, , [LLM Agents for Education, root [Pedagogical Agents, sa [Teaching Assistance §3.1, ta [Classroom Simulation (CS), ta [EduAgent (Xu et al., 2024a); TeachTune (Jin et al., 2024b); CGMI (Jinxin et al., 2023) Classroom Simulacra (Xu et al., 2025);  Li et al. (2025a);  Bhowmik et al. (2024);  Zheng et al. (2025);  Hu et al. (2025b), ta_work ] ] [Feedback Comment Generation (FCG), ta [ PROF (Nair et al., 2024); SEFL (Zhang et al., 2025a); FreeText (Matelsky et al., 2023); AAAR-1.0 (Lou et al., 2024);  Guo et al. (2024a);  Estévez-Ayres et al. (2024);  Du et al. (2024) ,ta_work ] ] [Learning Resource
Recommendation (LRR), ta [Zaiane (2002); SKarREC (Li et al., 2024d);  Moon et al. (2024);  Abu-Rasheed et al. (2024b);  Abu-Rasheed et al. (2024a) ,ta_work ] ] ] [Student Support §3.2, ss [Adaptive Learning (AL), ss [EduAgent (Xu et al., 2024a); ChatTutor (Chen et al., 2024b); Park et al. (2024); Wang et al. (2025b); Liu et al. (2024c); Sonlu et al. (2024); Liu et al. (2024a) ,ss_work ] ] [Knowledge Tracing (KT), ss [Yang et al. (2024a); Xu et al. (2025); Scarlatos et al. (2025) ,ss_work ] ] [Error Correction and Detection (ECD), ss [Ye et al. (2022); Li et al. (2024c); Xu et al. (2024b); Repairagent (Bouzenia et al., 2024); Wang et al. (2025b); ErrorRadar (Yan et al., 2024c); CoT Rerailer (Wan et al., 2024) ,ss_work ] ] ] ] [Domain-Specific Educational Agents, ca [Science Learning §4.1, sl [Mathematics, sl [MathAgent (Yan et al., 2025a); TORA (Gou et al., 2023); MathChat (Wu et al., 2023); MACM (Lei et al., 2025); Xiong et al. (2024), sl_work ] ] [Physics, sl [NEWTON (Wang et al., 2023); Physics Reasoner (Pang et al., 2024); SGA (Ma et al., 2024c), sl_work ] ] [Chemistry, sl [ChemCrow (M. Bran et al., 2024); ChemAgent (Yu et al., 2024); ChemReasoner (Sprueill et al., 2024); Curie (Kon et al., 2025), sl_work ] ] [Biology, sl [ProtChat (Huang et al., 2024b); ProtAgents (Ghafarollahi and Buehler, 2024a); TourSynbio-Agent (Shen et al., 2024), sl_work ] ] [General Scientific Discovery, sl [PaperQA (Lála et al., 2023); PaperQA2 (Skarlinski et al., 2024); OpenScholar (Asai et al., 2024); SciAgents (Ghafarollahi and Buehler, 2024b); SciAgent (Ma et al., 2024a); SciToolAgent (Chen et al., 2025b); TheoremExplainAgent (Ku et al., 2025); LLM-SR (Shojaee et al., 2024); AI Scientist (Lu et al., 2024); ResearchAgent (Baek et al., 2024); Agent Laboratory (Schmidgall et al., 2025); DiscoveryWorld (Jansen et al., 2024), sl_work ] ] ] [Language Learning §4.2, ll [Reading, ll [ExpertEase (Mo and Hu, 2024); AgentSimp (Fang et al., 2025); Chen and Leitch (2024), ll_work ] ] [Writing, ll [Weaver (Wang et al., 2024b); CAELF (Hong et al., 2024); EvaAI (Lagakis and Demetriadis, 2024); Debate-to-Write (Hu et al., 2025a), ll_work ] ] [Translation, ll [Agent-SimT (Guo et al., 2024b); TransAgents (Wu et al., 2024), ll_work ] ], [Storytelling, ll [STARie (Li and Xu, 2023); StoryAgent (Sohn et al., 2024); LLaMS (Zhang et al., 2025c), ll_work ] ], [Speaking, ll [ELLMA-T (Pan et al., 2024); Spoken-WOZ (Si et al., 2023); FurChat (Cherakara et al., 2023), ll_work ] ] ] [Professional
Development §4.3, pd [Medical Education, pd [Talk2Care (Yang et al., 2024d); openCHA (Abbasian et al., 2023); MEDCO (Wei et al., 2024); Abd-Alrazaq et al. (2023); Steenstra et al. (2024); Agent Hospital (Li et al., 2024b), pd_work ] ] [Computer Science Education, pd [CodeAgent (Zhang et al., 2024b); ToolCoder (Ding et al., 2025); SWE-agent (Yang et al., 2025); AgentCoder (Huang et al., 2023c); MapCoder (Islam et al., 2024); Iris (Bassner et al., 2024); Chen et al. (2024d); HypoCompass (Ma et al., 2024b); AlgoBo (Jin et al., 2024a); Magentic-One (Fourney et al., 2024); AutoKaggle (Li et al., 2024e); AgentK v1.0 (Grosnit et al., 2024); SWE-Search (Antoniades et al., 2024); AgentLess (Xia et al., 2024); TRAVER (Wang et al., 2025a), pd_work ] ] [Law Education, pd [Lawbench (Fei et al., 2023); LegalBench (Guha et al., 2023); LegalAgentBench (Li et al., 2024a); DeliLaw (Xie et al., 2024a); Agentcourt (Chen et al., 2024c); Agentscourt (He et al., 2024a); AgentsBench (Jiang and Yang, 2024); LawLuo (Sun et al., 2024) , pd_work ] ] ] ] ]
Figure 2. Taxonomy of representative research on education agents.
3. Pedagogical Agents
Pedagogical agents are AI-driven systems designed to enhance both teaching and learning through automation and personalization. These agents can be broadly categorized into two types: Agents for Teaching Assistance, which automate key tasks for educators, and Agents for Student Support, which provide personalized learning guidance to students. By streamlining tasks for teachers and offering real-time, adaptive feedback to students, these agents improve instructional effectiveness, reduce workloads, and foster engaging, tailored learning experiences.
3.1. Agent for Teaching Assistance
Agents for teaching assistance are designed to support educators in the learning environment. These agents leverage LLMs to provide personalized, scalable, and efficient support across various aspects of the educational process. Their primary objectives are to enhance teaching quality, enrich student learning experiences, and reduce educators’ workload. By incorporating advanced agent capabilities—such as memory management for retaining contextual information, tool integration for accessing external databases and APIs, and planning for precise student profile modeling—these agents can effectively assist in key areas, including classroom simulation3.1.1), feedback comment generation3.1.2), and learning resource recommendation3.1.3).
3.1.1. Classroom Simulation (CS)
Classroom simulation refers to the ability of teaching agents to replicate and model various classroom scenarios, such as student-teacher dialogues, collaborative learning activities, and problem-solving tasks. These simulations create dynamic and interactive learning environments where educators can experiment with different teaching strategies, assess student reactions, and receive real-time feedback on how various pedagogical approaches may unfold. By simulating these classroom dynamics, educators can refine their methods, anticipate student challenges, and enhance overall instructional effectiveness, all without the constraints of a physical classroom setting.
Effective classroom simulation depends on accurately modeling student behavior. Previous studies (Xu et al., 2024a; Jin et al., 2024b; Li et al., 2025a) demonstrate that LLM-based agents can predict fine-grained student behaviors across diverse personas and past learning patterns, aligning closely with real teachers’ expectations. To enhance simulation, the CGMI framework (Jinxin et al., 2023) uses a tree-based cognitive architecture with memory, reflection, and planning modules to simulate roles like teacher, student, and supervisor, improving realism. Similarly, Classroom Simulacra (Xu et al., 2025) incorporates a transferable iterative reflection module for more accurate behavior simulation. These systems enable automated interactions that reduce educators’ task loads while broadening the exploration of student profiles. Simulations can also test educational strategies tailored to different profiles, enhancing teaching quality, as shown in studies by Bhowmik et al. (2024) and Zheng et al. (2025). Additionally,  Hu et al. (2025b) demonstrate how LLMs can refine teaching plans through integrated simulations. To sum up, classroom simulations can be leveraged to test different educational strategies tailored to diverse student profiles, ultimately enhancing the quality of education.
3.1.2. Feedback Comment Generation (FCG)
Providing timely, relevant, and constructive feedback is a cornerstone of effective education. Teaching assistant agents can generate automated feedback comments on students’ assignments, quizzes, and projects. For example, Guo et al. (2024a) presents a training-free system for providing accurate feedback to students, which introduces a two-agent system. Specifically, Agent 1 generates initial feedback based on the students’ responses, while Agent 2 evaluates and refines this feedback to prevent over-praise and excessive inferences. Furthermore, Nair et al. (2024) designs a training strategy called PROF, which trains an automated LLM-based writing comment generator through reinforcement learning. This system adopts an iterative pipeline to simulate various student writing styles and incorporates a more advanced revision model (e.g., GPT-4) to provide the quality of the feedback as rewards. Similarly, SEFL (Zhang et al., 2025a) enhances feedback generation by having LLM agents role-play both students and teachers to generate data, which is then used to fine-tune models and improve feedback capabilities. These systems have also been deployed in real-world applications, such as FreeText (Matelsky et al., 2023), which pairs student responses with teacher-provided criteria, enabling the agent to identify strengths and weaknesses and provide targeted feedback for improvement. Beyond traditional feedback, advanced LLM agents are now capable of handling more complex, expertise-intensive tasks. For example, Du et al. (2024) explores the potential of LLM agents as assistants for natural language processing (NLP) paper reviewing tasks, while AAAR-1.0 (Lou et al., 2024) evaluates agents’ capabilities in areas such as equation inference, experiment design, paper weakness analysis, and review critique, revealing their potential in conducting sophisticated research tasks.
However, another line of research highlights that agent-generated feedback still faces challenges in handling complex tasks, such as programming and the review of professional academic papers (Estévez-Ayres et al., 2024; Lou et al., 2024). For instance, these agents may struggle with concepts like starvation and deadlocks, leading to inaccurate or incorrect feedback. Future work could focus on integrating external tools (e.g., search engines) and enhancing memory mechanisms to better address complex problem-solving scenarios, as well as refining the personalization of feedback across diverse learning contexts.
3.1.3. Learning Resource Recommendation (LRR)
To ensure that students access the most suitable learning resources aligned with their learning path and knowledge domain, it is crucial to develop an effective recommendation system. Zaiane (2002) first introduced recommendation systems into e-learning, utilizing web mining techniques to suggest online learning activities or shortcuts on course websites. Since then, the integration of LLM-based agents has opened new avenues for learning resource recommendations. These recommendations can be generated using both retrieval-based and generation-based methods. Retrieval-based methods involve agents accessing a database or their own memory to suggest existing resources—such as textbooks, research papers, or online content—based on student queries, past behavior, or content similarities (Shahzad et al., 2025; Li et al., 2024d). In contrast, generative methods create new learning content tailored to an individual student’s learning style, knowledge gaps, and interests (Moon et al., 2024). Moreover, to enhance the understanding and acceptance of recommended content, these agents need to provide reasons for their recommendations. Explaining the rationale behind suggestions fosters trust and enables students to make more informed decisions about the resources they engage with. For example, Abu-Rasheed et al. (2024a, b) incorporate knowledge graphs to recommend human-curated sources of information. This approach not only enhances the interpretability of recommendations but also reduces the risk of generating misinformation, thereby improving the quality of the learning experience.
Future work should focus on integrating agents with adaptive learning systems to dynamically adjust recommendations based on real-time student performance. Also, a hybrid approach combining generative and retrieval-based methods could enhance both accuracy and diversity. Furthermore, incorporating multi-modal content—including interactive media, video, and immersive simulations—could further enrich the learning experience.
3.2. Agent for Student Support
LLM agent-based student support systems enable independent, personalized learning by providing real-time academic assistance without requiring direct teacher involvement. Unlike traditional rule-based learning systems, they deliver interactive, tailored feedback, allowing students to learn at their own pace. Broadly, their core functionalities include adaptive learning3.2.1), knowledge tracing3.2.2), and error correction and detection3.2.3).
3.2.1. Adaptive Learning (AL)
LLM agents can automate tasks, making it possible to build self-sustaining adaptive learning systems without teacher involvement. These systems operate within an action-memory framework, continuously refining instruction based on student profiling, content selection, and difficulty adaptation. The agent maintains a structured student profile, dynamically selecting materials based on past interactions. After each task, it analyzes student behavior and adjusts teaching strategies accordingly, enhancing personalization and responsiveness.
Several implementations exemplify this adaptive approach. EduAgent (Xu et al., 2024a) introduces a structured profiling mechanism comprising four distinct cognitive patterns: gaze behavior linked to physiological memory, motor behavior mapped to motor memory, cognitive state associated with cognitive memory, and post-course assessments contributing to knowledge memory. This structured representation enhances adaptive decision-making by providing a multi-faceted view of student learning states. Chen et al. (2024b) propose a system consisting of interaction, reflection, and reaction, with each component composed of specific LLM tools and memory modules. Furthermore, a meta-agent is introduced to control the information flow through these agents.
Building on these adaptive mechanisms, LLM agents operate within a state-action framework, allowing them to dynamically adjust instructional strategies in response to a student’s evolving cognitive and affective state. The state space in these systems encodes the student’s cognitive and affective profile, while the action space governs how the agent adjusts its teaching approach. A typical cognitive state representation includes tracking a student’s knowledge proficiency, comprehension levels, and misconceptions (Park et al., 2024; Liu et al., 2024c, a), allowing the agent to tailor explanations, adjust difficulty levels, and reinforce concepts dynamically. Recent studies highlight the importance of affective state modeling, as emotional factors such as motivation, interest, and self-efficacy significantly influence learning outcomes. For instance, Park et al. (2024) propose an affective state model that enables agents to adjust feedback tone, provide encouragement, and regulate pacing to maintain engagement. Another crucial dimension of adaptation involves learning preferences and personality traits. These studies integrate personality with memory design, tracking the psychological state of the students. Wang et al. (2025b) integrate learning preferences into state modeling, recognizing that students process information differently depending on instructional format and modality. Adapting content to these preferences enhances retention and learning efficiency. Moreover, Liu et al. (2024c) apply the Big Five personality model (Roccas et al., 2002) to personalize tutoring strategies, acknowledging that individual differences shape learning experiences (Sonlu et al., 2024).
Recent approaches explore multi-agent systems for adaptive learning, where specialized agents collaborate to enhance personalization. For example, Wang et al. (2025b) design five agents—Gap Identifier, Learner Profiler, Dynamic Learner Simulator, Learning Path Scheduler, and Content Creator—to deliver goal-oriented, personalized instruction. Similarly, OATutor (Pardos et al., 2023) provides an experimental platform for modular adaptive learning, allowing researchers to design scalable, domain-general tutoring agents.
3.2.2. Knowledge Tracing (KT)
Knowledge tracing is a fundamental component of intelligent tutoring systems, enabling the tracking of a learner’s evolving understanding over time to predict future performance and knowledge retention. Traditional knowledge tracing methods often rely on statistical models or deep learning approaches to estimate student mastery levels. In contrast, LLM agents introduce a more dynamic and personalized approach, leveraging their ability to process natural language interactions, infer conceptual mastery, and adapt instructional strategies accordingly.
Recent advancements have explored multi-agent frameworks for knowledge tracing. For instance, Yang et al. (Yang et al., 2024a) propose a multi-agent system with three specialized agent roles: administrator, judger, and critic. In this framework, the administrator delegates knowledge tracing tasks to judgers, who collaborate through discussions to assess the student’s cognitive state. The critic agent then evaluates the outcome and determines whether the assessment criteria are met, ensuring a structured yet flexible knowledge tracing process. Other agent-based approaches explore alternative strategies for modeling student knowledge. Xu et al. (2024a) propose simulating students as different personas, allowing agents to adaptively trace knowledge progression based on varied learning profiles. Meanwhile, Scarlatos et al. (2025) employ dialogue-driven interactions to probe students’ conceptual boundaries, using conversational exchanges to refine knowledge estimation dynamically.
3.2.3. Error Correction and Detection (ECD)
Error detection and correction are critical components of intelligent tutoring systems, enabling students to refine their understanding by receiving real-time, context-aware feedback. LLM agents can identify errors in student responses across various domains, including academic writing, programming, and mathematical reasoning. By dynamically adapting feedback to the learner’s proficiency level, these agents serve as intelligent reviewers, debuggers, and writing assistants (Ye et al., 2022; Li et al., 2024c).
Agent-based systems leverage state representations and adaptive inference mechanisms to track error patterns and misconceptions dynamically (Park et al., 2024; Liu et al., 2024c; Wang et al., 2025b; Bouzenia et al., 2024). Recent advancements extend this capability into the multi-modal domain, incorporating direct analysis of student-generated drafts. Xu et al. (2024b) propose a multimodal LLM framework that processes handwritten or digitally drafted student work. The system first extracts and converts draft content into natural language, enabling the LLM-based agent to interpret and analyze handwritten responses. The agent then provides indirect yet effective instructional feedback, guiding students toward self-correction and deeper comprehension. Moreover, CoT Rerailer (Wan et al., 2024) designs a derailment identification process and a rerailment process to conduct error detection. Zhang et al. (2025b) propose the MathCCS benchmark and develop a sequence error analysis framework, adopting multi-agent collaboration. As the first benchmark for multimodal error detection, ErrorRadar (Yan et al., 2024c) can provide a data foundation for multimodal agents in error detection task.
4. Domain-Specific Educational Agents
Recent research on LLM agents in education has also shown growing interest in domain-specific applications. We explore their use in science learning, language learning, and professional development, focusing on their algorithmic frameworks, agentic designs, and relevant datasets and benchmarks.
4.1. Agent for Science Learning
An agent for science learning is an intelligent system powered by LLMs, designed to assist students in acquiring and applying scientific knowledge through personalized, interactive experiences (Yan et al., 2025b; Raihan et al., 2025; Ng et al., 2024; Brown and Cruz Castro, 2025). The significance of these agents in education lies in their ability to offer tailored feedback, enhance conceptual understanding, and promote active engagement with complex scientific ideas. In the following sections, we explore the impact of LLM agents in four key scientific disciplines: mathematics4.1.1), physics4.1.2), chemistry4.1.3), and biology4.1.4), as well as their broader contributions to general scientific discovery4.1.5).
4.1.1. Mathematics
In mathematics, LLM agents provide substantial support by helping students navigate complex problems and reinforcing their understanding of abstract concepts (Yan et al., 2024b; Xiong et al., 2024; Swan et al., 2023; Yan et al., 2024c; Wu, 2025; Mitra et al., 2024). For instance, Gou et al. (2023) introduce TORA (Tool-integrated Reasoning Agents), a framework that integrates natural language reasoning and program-based tool use to handle mathematical reasoning. MathAgent (Yan et al., 2025a) similarly proposes Mixture-of-Math-Agent framework to address multimodal error detection in real-world K-12 scenarios, and flexibly transform the visual information of different types of questions into forms that are more easily understood by LLMs (e.g., converting plane geometry images into formalized expression). Additionally, MathChat (Wu et al., 2023) serves as a conversational mathematical problem-solving agent, which consists of a chat-based LLM agent and a tool-based user agent. Furthermore, Xiong et al. (2024) propose to use reinforcement learning from human feedback (RLHF) to further improve tool-integrated agents for mathematical problem-solving, and formulate this method as a Markov decision process, distinguishing it from the typical contextual bandit approach used in RLHF. Besides, MACM (Lei et al., 2025) discuss the limitations of LLMs in handling complex mathematical logical deduction, thus introducing a multi-agent system, which comprises three interactive agents: Thinker, Judge, and Executor.
4.1.2. Physics
In the field of physics, LLM agents help students make sense of challenging concepts and offer interactive tools to simulate physical phenomena (Pang et al., 2024; Mower and Bou-Ammar, 2025; Barman et al., 2025; Feng et al., [n. d.]; Jiang and Jiang, 2024; Yan and Lee, 2024). Wang et al. (2023) introduce NEWTON, the first pipeline and benchmark to explore the physical reasoning abilities of LLMs. Furthermore, Kortemeyer (2023) describe a case study exploring if an LLM agent can pass an introductory calculus-based physics course. In addition, Physics Reasoner (Pang et al., 2024), a novel knowledge-augmented framework for physics problem-solving, leverages a comprehensive formula set and detailed checklists to ensure accuracy and completeness. It can serve as an agent consisting of three stages - problem analysis, formula retrieval, and guided reasoning. Besides, Ma et al. (2024c) describe the Scientific Generative Agent (SGA), a bilevel optimization framework designed for physical scientific discovery, and highlight the use of LLMs for generating and revising scientific hypotheses and implementing an exploit-and-explore strategy.
4.1.3. Chemistry
Chemistry education also benefits greatly from LLM agents, which can explain molecular structures, chemical reactions, and experimental processes in an engaging and interactive way (Ramos et al., 2025; M. Bran et al., 2024; Yu et al., 2024; Guo et al., 2023; Tsai et al., 2023). For example, ChemCrow (M. Bran et al., 2024) is the first LLM chemistry agent capable of autonomous planning and execution of chemical syntheses, including an insect repellent and three organocatalysts. Yu et al. (2024) further present ChemAgent, an enhanced chemistry agent improved over ChemCrow, with a focus on two essential cognitive abilities of chemistry problem-solving: reasoning and grounding. Besides, Curie (Kon et al., 2025) is an agent framework aimed at incorporating rigor into the experimentation process via three core elements: an intra-agent rigor module to boost reliability, an inter-agent rigor module to ensure systematic control, and an experiment knowledge module to improve interoperability. Recent studies have explored the capabilities of LLMs in complex chemical discovery (Yang et al., 2024c; Ruan et al., 2024; Sprueill et al., 2024; Moret et al., 2023; Jablonka et al., 2024), and their potential can be advanced by leveraging the interactivity of agent-based tool use and the flexibility of planning strategies (Song et al., 2024; Ramos et al., 2025).
4.1.4. Biology
In biology, LLM agents enhance learning by offering detailed explanations of biological processes and providing interactive experiences to explore living systems (Yan et al., 2025b; Bhattacharya et al., 2024; Sripathi et al., 2024; Zhao et al., 2025; Gao et al., 2024). For example, ProtChat (Huang et al., 2024b) is a multi-agent tool leveraging GPT-4 and Protein Language Models for seamless protein analysis automation, thus evolutionizing the complexities of protein sequence interpretation. ProtAgents (Ghafarollahi and Buehler, 2024a) is introduced as a multi-agent modeling framework that combines state-of-the-art LLMs with diverse tools to tackle protein design and analysis. It consists of a team of agents: User, Planner, Assistant, Critic, and Group Chat Manager. Besides, Shen et al. (2024) present TourSynbio-Agent, an innovative agent framework that leverages TourSynbio-7B’s protein understanding ability to perform various protein engineering tasks, such as mutation analysis, inverse folding, and visualization.
4.1.5. General Scientific Discovery
LLM agents support general scientific discovery by assisting students in data interpretation, hypothesis testing, and creative problem-solving (Yan et al., 2025b; Chen et al., 2024a; Ghafarollahi and Buehler, 2024b; Chen et al., 2025b; Schmidgall et al., 2025). These LLM agents, such as PaperQA (Lála et al., 2023), PaperQA2 (Skarlinski et al., 2024), OpenScholar (Asai et al., 2024), SciAgents (Ghafarollahi and Buehler, 2024b), TheoremExplainAgent (Ku et al., 2025), and LLM-SR (Shojaee et al., 2024), can analyze complex scientific datasets, helping students uncover patterns and trends that may not be immediately apparent. In addition, Narayanan et al. (2024) present Aviary, an extensible gymnasium for language agents for three challenging scientific tasks: manipulating DNA constructs for molecular cloning, answering research questions by accessing scientific literature, and engineering protein stability. Furthermore, both SciAgent (Ma et al., 2024a) and SciToolAgent (Chen et al., 2025b) extend to a tool-augmented scientific reasoning setting with the help of domain-specific tools. Besides, Agent Laboratory (Schmidgall et al., 2025) emerges as an agent framework that automates the research process of three phases (Literature Review, Experimentation, and Report Writing) via various LLM agents (PhD, Postdoc, ML Engineer, etc.).
4.2. Agent for Language Learning
The integration of LLM agents into language learning is revolutionizing how core competencies—reading, writing, listening, and speaking—are taught and practiced (Ye et al., 2025b). These skills form the foundation of effective communication and language acquisition, and recent advancements in LLM-based agents have significantly enhanced how learners interact with and acquire these skills. Below, we introduce recent advancements in each subdomain, highlighting the role of LLM agents in enhancing pedagogical outcomes in language acquisition for students and second language (L2) speakers through engaging and adaptive approaches (Huang et al., 2023b; Ye et al., 2023a, b).
4.2.1. Reading
Reading comprehension is a vital component of language learning, and LLM agents are playing an increasingly important role in enhancing students’ reading abilities. For instance, ExpertEase (Mo and Hu, 2024) employs a multi-agent framework to adapt documents for grade-specific audiences, simulating expert-teacher-student collaboration to enhance comprehension. AgentSimp (Fang et al., 2025) tackles document-level simplification by leveraging multiple agents with distinct roles to ensure coherence and accessibility. Additionally, various LLMs (Adetayo et al., 2024) have been used as academic reading companions, demonstrating improved engagement and understanding of complex qualitative texts in educational settings (Chen and Leitch, 2024).
4.2.2. Writing
The development of writing skills has benefited significantly from NLP tasks like explainable grammatical error correction (EXGEC) (Ye et al., 2024b, 2025a; Zou et al., 2025) and automatic essay scoring (AES) systems (Su et al., 2025). Weaver (Wang et al., 2024b), a family of LLMs fine-tuned for writing tasks (Ye et al., 2023c, 2024c), outperforms generalist LLMs like GPT-4 in generating human-like narratives. Moreover, Weaver natively supports retrieval-augmented generation (RAG) and function calling, serving as a qualified foundational model for LLM agents. For interactive feedback on student essays, CAELF (Hong et al., 2024) introduces a multi-agent framework that enables interactive essay feedback. By combining Teaching-Assistant agents’ evaluations with teacher-agent arbitration, students can contest grades and engage with the feedback, addressing the “black box” limitations of traditional automated scoring. Inspired by the process of human debate, Debate-to-Write (Hu et al., 2025a) construct a persona-based multi-agent framework that can enable agents to collaboratively debate, discuss ideas, and form a comprehensive plan for argument writing.
4.2.3. Translation
LLM agents demonstrate remarkable advancements in both simultaneous (Koshkin et al., 2024; Guo et al., 2024c) and literary translation (Cheng et al., 2024; Škobo and Petričević, 2023). Translation tasks benefit from LLM agents through their ability to integrate specialized tools and orchestrate multi-agent collaboration. Agent-SiMT (Guo et al., 2024b) combines the decision-making capabilities of a Simultaneous Machine Translation (SiMT) policy agent with the generative power of a translation agent, achieving state-of-the-art performance in simultaneous translation by dynamically balancing reading and generation actions. For literary translation, TransAgents (Wu et al., 2024) employs a multi-agent framework to replicate the complex workflows of human translation teams, addressing cultural nuances and stylistic challenges through collaborative reasoning. This approach not only improves translation quality but also extends LLM applications to linguistically and culturally rich domains. These contributions underscore the importance of tool use and agent collaboration in advancing translation education (Zheng et al., 2024).
4.2.4. Storytelling
Storytelling applications leverage LLM agents to create immersive and interactive learning experiences (Simon and Muise, 2022). STARie (Li and Xu, 2023), a peer-like embodied conversational agent, integrates multimodal tools such as speech synthesis and facial animation to scaffold children’s storytelling, fostering narrative creativity and oral communication skills (Beredo and Ong, 2021; Cassell, 2022). StoryAgent (Sohn et al., 2024) combines top-down story drafting with bottom-up asset generation to transform simple prompts into coherent, multi-modal digital narratives. By automating complex storytelling workflows (Liem et al., 2023), it democratizes content creation and enhances engagement in language learning. LLaMS (Zhang et al., 2025c), a multi-modal agent framework, is designed to generate multi-modal human-level stories characterized by expressiveness and consistency, incorporating the Story-Adapter module for long image sequence illustration. These systems demonstrate the potential of LLM agents to support both cognitive and creative aspects of language education for children.
4.2.5. Speaking
LLM agents are revolutionizing spoken language education by integrating reasoning and multi-agent collaboration to build adaptive dialogue systems (Liu et al., 2024b; Balan et al., 2024). ELLMA-T (Pan et al., 2024) employs contextual reasoning and role-playing in social VR environments to provide personalized feedback and language assessments, enabling learners to practice speaking in realistic scenarios (Lim et al., 2024; Li et al., 2025b). SpokenWOZ (Si et al., 2023) introduces a large-scale benchmark for task-oriented spoken dialogue, highlighting the importance of reasoning and multi-turn interaction in addressing real-world conversational challenges. FurChat (Cherakara et al., 2023), an embodied conversational agent, combines verbal and non-verbal communication cues to simulate natural interactions, making it a valuable tool for improving speaking skills through immersive and realistic practice. By employing multimodal signals such as speech and gestures, SpeechAgents (Zhang et al., 2024c) enhances the authenticity of dialogue simulations, capturing consistent content, natural rhythm, and rich emotional expression. Through Multi-Agent Tuning (Liang et al., 2024), it optimizes LLM capabilities for large-scale simulations involving up to 25 agents, enabling applications like drama creation and audio novel generation.
4.3. Agent for Professional Development
Agents for professional development harness the capabilities of LLMs to offer scalable, adaptive, and context-aware learning experiences tailored to domain-specific needs. This section summarizes how recent studies develop agents to revolutionize professional training in fields including medical4.3.1), computer science4.3.2), and law education4.3.3).
4.3.1. Medical Education
The deployment of LLM agents in healthcare has created new opportunities for personalized, interactive, and scalable systems, with several health agents introduced (Shusterman et al., 2025) such as Talk2Care (Yang et al., 2024d) and openCHA (Abbasian et al., 2023). Additionally, Abd-Alrazaq et al. (2023) highlight the educational potentials of LLMs in crafting personalized curricula, adaptive learning plans, and dynamic assessment tools for medical education, while concurrently addressing challenges including algorithmic bias, misinformation, and privacy issues. MEDCO (Wei et al., 2024), a multi-agent system, has the capacity to replicate real-world medical training environments through agent collaboration with virtual patients, expert physicians, and radiologists, enhancing interdisciplinary learning and peer interaction. Furthermore, Abbasian et al. (2023) introduce openCHA as a personalized LLM-powered framework that integrates external resources and orchestrates multi-step problem-solving for complex healthcare queries (Ye et al., 2024a), emphasizing tool use and action planning. Beyond traditional education, Steenstra et al. (2024) explore LLMs in creating fantasy narrative games for adolescent health education, demonstrating the agents’ ability to generate engaging, doctor-validated content that enhances knowledge retention through gamification. Li et al. (2024b) present Agent Hospital, a simulation environment where LLM-driven agents evolve through autonomous interactions, demonstrating significant improvements in medical reasoning and performance on benchmarks like MedQA (Jin et al., 2021) after treating thousands of simulated patients. Collectively, these investigations highlight the versatility of LLM agents within medical education, demonstrating their abilities in reasoning, collaboration, tool integration, and adaptive learning to effectively address a broad spectrum of educational and clinical challenges (Karabacak and Margetis, 2023; Tian et al., 2024b; Ullah et al., 2024).
4.3.2. Computer Science Education
An agent for computer science (CS) education greatly enhances learning by providing personalized guidance on coding, debugging, and understanding CS principles (Ma et al., 2024b; Lee and Song, 2024; Kosar et al., 2024; Liu et al., 2024d). For example, CodeAgent (Zhang et al., 2024b) serves as an LLM agent framework for repo-level code generation, incorporating external tools such as WebSearch and DocSearch. Recent studies have demonstrated the potential of agent-based code generation systems such as ToolCoder (Ding et al., 2025), SWE-agent (Yang et al., 2025), AgentCoder (Huang et al., 2023c), and MapCoder (Islam et al., 2024), which can significantly enhance students’ coding efficiency (Jin et al., 2024a; Wang et al., 2025a; Frankford et al., 2024). Furthermore, Bassner et al. (2024) introduce Iris, an LLM-driven virtual tutor designed to offer personalized, context-aware assistance to CS students within the interactive learning platform Artemis. Besides, Chen et al. (2024d) propose Learning-by-Teaching (LBT) as an effective pedagogical strategy for CS education, and leverage the advantages of LLM agents (e.g., contextual conversation & learning from demonstrations).
4.3.3. Law Education
LLM agents leverage pre-trained legal knowledge, interactive capabilities, and reasoning skills to support law education through judicial interpretation, moot court simulation, and case analysis (Chen et al., 2024c; Nelson, 2024; Lai et al., 2024; Yuan et al., 2024). However, evaluations from LawBench (Fei et al., 2023) and LegalBench (Guha et al., 2023) reveal that LLMs struggle with legal knowledge application and judicial aid. LegalAgentBench (Li et al., 2024a) further highlights their limitations in multi-hop reasoning and defense statement writing, showing that LLM agents require significant improvements to effectively assist in complex legal tasks. Despite these challenges, LLM agents are emerging as valuable tools for moot court simulations, a crucial component of legal reasoning and advocacy training. DeliLaw (Xie et al., 2024a) enhances law education by integrating legal and case retrieval modules, enabling students to practice legal research, case analysis, statutory interpretation, and mock consultations. LawLuo (Sun et al., 2024) applies a multi-agent framework with retrieval-augmented generation to simulate multi-turn legal consultations, improving personalization and ambiguity handling. Similarly, AgentCourt (Chen et al., 2024c) and AgentsCourt (He et al., 2024a) simulate courtroom interactions and judicial decision-making, providing a realistic training ground for law students. AgentsBench (Jiang and Yang, 2024) extends this by offering multi-agent legal reasoning and case analysis, further advancing AI-driven legal education.
5. Challenges & Future Directions
In this section, we discuss key challenges that must be addressed to ensure the effectiveness, reliability, and ethical deployment of LLM agents for education. We focus on three critical areas: privacy, bias, and fairness concerns; hallucination and overreliance; and integration with existing educational ecosystems. For each challenge, we outline potential research directions to enhance the robustness and applicability of LLM agents in learning environments.
5.1. Privacy, Bias and Fairness Considerations
The integration of LLM agents into educational applications offers transformative potential but also raises significant ethical concerns, particularly regarding privacy and bias.
LLM agents process vast datasets, often containing sensitive personal information, leading to potential privacy risks. Studies highlight low technological readiness and insufficient privacy measures in educational contexts (Yan et al., 2024a). Emerging research (He et al., 2024b; Gan et al., 2024; Zhang et al., 2024d; Hua et al., 2024; Huo et al., 2025; Chen et al., 2025a) underscores new privacy and security concerns, emphasizing the need for stronger data protection mechanisms (Huang, 2023; Ismail, 2025; Khan, 2024). Additionally, bias in LLMs remains a pressing concern, as models trained on large datasets can inadvertently reinforce stereotypes and disparities, affecting educational fairness. Recent work calls for bias mitigation strategies to promote equitable learning experiences (Adewumi et al., 2024; Aird et al., 2024; Mehrotra et al., 2024). Addressing these biases is essential to ensuring inclusive, unbiased educational outcomes.
To overcome the above issues, a number of future directions can be explored: (i) Privacy-preserving memory management: develop context-aware memory mechanisms that allow agents to retain useful learning progress while forgetting sensitive user data when necessary. (ii) Bias detection and mitigation: develop automated fairness-checking models that evaluate real-time content generated by LLM agents to detect biased explanations, language, or examples. (iii) Culturally adaptive LLM Agents for global education: train multilingual, culturally aware educational agents that dynamically adjust explanations based on regional learning norms, historical perspectives, and diverse curricula.
5.2. Hallucination and Overreliance
The phenomenon of ”hallucinations” (Zhang et al., 2023), where LLMs generate plausible-sounding but incorrect or nonsensical information, poses challenges to their reliability in educational contexts.
In educational settings, hallucinations can mislead learners by presenting plausible but inaccurate information as factual, leading to misconceptions (da Silva et al., 2024; Jho, 2024; Ho et al., 2024). For example, AI-generated content may fabricate historical events or scientific data, which students may unknowingly accept as true. This issue is particularly concerning given the authoritative tone of AI-generated responses, making errors harder to detect and correct. Also, recent researches (Milano et al., 2023; Krupp et al., 2024; Adewumi et al., 2023) show that overreliance on AI-generated content may hinder genuine skill acquisition and impede in-depth learning.
Some directions can be explored to mitigate hallucinations in LLM agents for education: (i) Self-correcting AI tutors: develop LLM agents with self-reflection capabilities, where models review, verify, and refine their own generated content before presenting it to students. (ii) Hybrid Human-AI feedback loops for educational content verification: develop teacher-in-the-loop AI systems where educators can review and correct AI-generated responses, refining LLM performance over time. (iii) Pedagogical-aware educational agents: design agentic frameworks that align with human expert pedagogical practices.
5.3. Integration of LLM agents into Existing Educational Ecosystems
The integration of LLM agents into education presents opportunities for personalized learning, curriculum development, and project-based learning (PBL), but also raises challenges related to structured deployment, equitable access, and privacy.
One major challenge is the lack of structured frameworks for integrating LLM agents into educational systems. While models like the FOKE framework (Hu and Wang, 2024) combine foundation models, knowledge graphs, and prompt engineering to provide interactive and explainable learning services, broader adoption requires scalable models that can be validated in diverse real-world educational settings. Additionally, LLMs have been explored as tools to enhance creativity and collaboration in PBL, supporting students through brainstorming, problem-solving, and project execution. However, studies indicate that their effectiveness is limited by the absence of structured guidance frameworks that help educators and students seamlessly incorporate LLM agents into PBL workflows (Zha et al., 2024). Another critical challenge is ensuring equitable access to LLM-powered educational tools, particularly in underfunded schools and institutions with limited AI infrastructure. Platforms such as AI-VERDE (Mithun et al., 2025) aim to democratize access by providing a unified LLM-as-a-platform service with built-in access control, privacy-preserving mechanisms, and budget management. However, achieving widespread adoption still depends on scalable and cost-effective deployment strategies that can support educational institutions at different resource levels.
Future research should focus on developing standardized frameworks to guide the structured deployment of LLMs in personalized learning, PBL, and assessment. Expanding models like FOKE with adaptive learning strategies, multimodal content, and real-time feedback could enhance instructional effectiveness. Additionally, integrating interactive AI tutors that support student collaboration, project tracking, and contextual guidance would further improve PBL applications. Ensuring cost-effective AI deployment through cloud-based and decentralized models would make LLM-powered learning tools more accessible to a wider range of institutions. Furthermore, providing educators with AI literacy training and oversight tools is essential for responsible and effective integration.
6. Conclusion
In this survey, we provided a comprehensive review of LLM agents for education, highlighting their potential to revolutionize personalized learning, intelligent tutoring, and pedagogical automation. We introduced a task-centric taxonomy, categorizing LLM agents into Pedagogical Agents and Domain-Specific Educational Agents, and analyzed their applications across various educational domains. Furthermore, we discussed key challenges and future research directions, including ethical considerations, bias mitigation, and real-world integration. To support further advancements, we compiled essential datasets and benchmarks to facilitate research in this emerging field. As LLM agents continue to evolve, their impact on engagement, automation, and personalized learning will grow, but ensuring their effectiveness, reliability, and ethical deployment remains a key challenge. We hope this survey serves as a foundation for future research, driving innovations in AI-driven education and advancing educational equity.
Appendix A Datasets & Benchmarks
In Table 1, we provide a comprehensive summary of publicly available datasets and benchmarks designed to evaluate LLM agents for education across various domains. It categorizes resources based on their primary goal, target users, subject domain, education level, language, modality and dataset size. We hope this collection can support and advance research on LLM agents for education.
Several datasets are designed to evaluate the pedagogical agents, such as ASSIST09 (Feng et al., 2009) and Junyi (Chang et al., 2015), which support knowledge tracing (KT) in K-12 math education, while others like EduAgent (Xu et al., 2024a) facilitate adaptive learning (AL) by dynamically adjusting content based on student profiles. In addition, error correction and detection (ECD) datasets, such as Virtual Teacher (Xu et al., 2024b) and MathCCS (Zhang et al., 2025b), assess LLM agents’ ability to identify and rectify student mistakes in math learning. Other datasets cater to writing, reading, and language learning, including FABRIC (Han et al., 2023), EssayJudge (Su et al., 2025), and EXCGEC (Ye et al., 2024b), which focus on feedback and generation (FCG) for student essays. MultiSim (Ryan et al., 2023) and Wang et al. (2024c) provide multi-lingual translation and storytelling benchmarks, expanding LLM capabilities beyond English-language education.
Several datasets support domain-specific educational agents across science, law, medicine, and computer science. Beyond their primary goal of evaluating pedagogical ability, these datasets assess LLM agents in domain-specific applications. They provide insights into how LLM agents can be adapted for specialized instruction, evaluating their ability to deliver subject-specific knowledge, facilitate problem-solving, and enhance interactive learning experiences across diverse educational fields. ScienceAgentBench (Chen et al., 2024a) and TheoremExplainBench (Ku et al., 2025) assess scientific reasoning and theorem explanation, while ML-Bench (Tang et al., 2023) and MLAgentBench (Huang et al., 2023a) focus on machine learning education. In law education, datasets like LawBench (Fei et al., 2023), LegalBench (Guha et al., 2023), and AgentCourt (Chen et al., 2024c) evaluate legal knowledge application, case analysis, and court simulations. Medical education datasets, including MedBench (Cai et al., 2024) and OmniMedVQA (Hu et al., 2024), test clinical reasoning and medical knowledge retrieval. For computer science, SWE-Bench (Yang et al., 2025) and Programming Feedback (Estévez-Ayres et al., 2024) assess code generation, debugging, and software engineering instruction. These benchmarks help refine LLM agents for specialized tutoring, enhancing AI-driven learning in professional fields.
Table 1. Summary of existing datasets and benchmarks of LLM agents for education.
Dataset&Benchmark Goal User Domain Level Language Modality Amount Source
ASSIST09 KT Student Math K12 EN text 227k (Feng et al., 2009)
Junyi KT Student Math K12 ZH text 2.5M (Chang et al., 2015)
EduAgent AL & CS Student - Graduate EN text & image 1,015 (Xu et al., 2024a)
MathDial AL & KT Student Math K12 EN text 45 (Macina et al., 2023)
MultiArith AL Student Math K12 EN text 180 (Xu et al., 2024a)
CoMTA KT Student Math K12 EN text 153 (Scarlatos et al., 2025)
MaCKT KT Student Math K12 EN text 452 (Yang et al., 2024a)
Virtual Teacher ECD Student Math K12 ZH text & image 420 (Xu et al., 2024b)
MathCCS ECD Student Math K12 ZH text & image 420 (Zhang et al., 2025b)
MathTutorBench ECD & FCG Student Math K12 EN text 7 tasks (Macina et al., 2025)
MultiSim - Student Reading - Multi-lingual text 1.7M (Ryan et al., 2023)
FABRIC FCG Student Writing - EN text 1,782 (Han et al., 2023)
EssayJudge FCG Student Writing - EN text & image 1,054 (Su et al., 2025)
PROF FCG Teacher Writing - EN text 363 (Nair et al., 2024)
EXCGEC FCG Student Writing - ZH text 8,216 (Ye et al., 2024b)
Wang et al. (2024c) - All Translation - Multi-lingual text 70K (Wang et al., 2024c)
NewEpisode - Student Storytelling - EN text & image 24.5K (Wang et al., 2024e)
SD-Eval ECD Student Speaking - EN speech 7,303 (Ao et al., 2024)
Programming Feedback FCG Teacher Computer Science - Code text 52 (Estévez-Ayres et al., 2024)
Review Critique FCG All Computer Science - EN text 440 (Du et al., 2024)
AAAR-1.0 FCG All Computer Science - EN text & image 1,000 (Lou et al., 2024)
ScienceAgentBench - All General Science - EN text 102 (Chen et al., 2024a)
TheoremExplainBench - All General Science - EN text 240 (Ku et al., 2025)
MLGym-Bench - All General Science - EN text 13 tasks (Nathani et al., 2025)
ML-Bench - All Machine Learning - EN text & image 9,641 (Tang et al., 2023)
MLAgentBench - All Machine Learning - EN text 13 tasks (Huang et al., 2023a)
SciCode - All General Science - EN text 338 (Tian et al., 2024a)
BLADE - All General Science - EN text 12 tasks (Gu et al., 2024)
DiscoveryBench - All General Science - EN text 1,167 (Majumder et al., 2024)
SUPER - All General Science - EN text 796 (Bogin et al., 2024)
E-EVAL - All Math & Language & General Science K12 ZH text 4,351 (Hou et al., 2024)
MedBench - All Medical - ZH text 40,041 (Cai et al., 2024)
CMB - All Medical - ZH text 280,839 (Wang et al., 2024a)
OmniMedVQA - All Medical - ZH text & image 127,995 (Hu et al., 2024)
MedEval - All Medical - EN text & image 22,779 (He et al., 2023)
OSWorld - All Computer Science - EN text & image 369 (Xie et al., 2024b)
Spider2-V - All Computer Science - EN text & image 812 (Cao et al., 2024)
VisualWebArena - All Computer Science - EN text & image 910 (Koh et al., 2024)
WebArena - All Computer Science - EN text & image 812 (Zhou et al., 2023)
SWE-Bench - All Computer Science - EN text 2,294 (Yang et al., 2025)
SWE-Bench M - All Computer Science - EN text & image 617 (Yang et al., 2024b)
Magentic-One - All Computer Science - EN text & image 617 (Fourney et al., 2024)
Lawbench - All Law - ZH text 20 tasks (Fei et al., 2023)
LegalBench - All Law - EN text 162 tasks (Guha et al., 2023)
LegalAgentBench - All Law - ZH text 300 tasks (Li et al., 2024a)
Agentcourt - All Law - ZH text 550 (Chen et al., 2024c)
SimuCourt - All Law - ZH text 420 (He et al., 2024a)